Help

(0.0034 / 0.035) times 100 Percent ionization = 9.
Answer #1

0.3990 atm = Molarity x 0.08206 x 298.15

Molarity = 0.01631

MgCl2 yields 3 particles per molecule

NaCl yield 2 particles per molecule

MgCl2: Molar mass = 95.211 g

NaCl: Molar mass = 58.443 g

If sample was 100% MgCl2 molarity = 3 x (0.5 / 95.21 ) = 0.01575 Mol/liter of particles

If sample was 100% NaCl molarity = 2 x (0.5 / 58.55 ) = 0.01711 Mol/liter of particles

Let x = Percent MgCl2

0.01575 x + 0.01711 x (1-x) = 0.01631

0.01575 x +0.01711 - 0.01711 x = 0.01631

0.0008 = 0.00136 x

x = 0.588

x % = MgCl2 = 58.8%

MgCl2 mass percent = 58.8 %